OpenStudy (anonymous):

Simplify 4sqrt(6) / sqrt(30) by rationalizing the denominator.

OpenStudy (anonymous):

What does rationalizing mean?

OpenStudy (anonymous):

4sqrt(6)/sqrt(6*5) = 4/sqrt(5)

OpenStudy (anonymous):

rationalizing means you make the denominator a rational number. you do this by multiplying the equation by \[\sqrt(30\div \sqrt(30)\] (which equals one)

OpenStudy (radar):

\[4\sqrt{6}\over \sqrt{30}\] It means making the denominator an expression that is rational, ie expressed as a number or a fraction ......no radical

OpenStudy (radar):

in this case you multiply both the denominator and the numerator by\[\sqrt{30}\] this is actually multiplying the expression by 1 and does not alter the actual value, just the appearance.

OpenStudy (radar):

\[4\sqrt{6}\sqrt{30}\over \sqrt{30}\sqrt{30}\]

OpenStudy (radar):

\[4\sqrt{180}\over 30\]\[24\sqrt{5}\over30\]\[4\sqrt{5}\over 5\]

OpenStudy (anonymous):

And then it becomes 4 / sqrt(5). Okay, thank you for the explanation!

OpenStudy (radar):

Not really, it is what it is. You could say \[.8\sqrt{5}\] but\[4 \over \sqrt{5}\]denominator is not rationalized as it is an irrational sqrt5, in addition it is also incorrect value